OpenStudy (anonymous):

point L is located at (4,-3) and M is located at (-8,5). find the y value of the point that lies halfway between L and M. A)1.5 b)4 c)3.5 d)1

OpenStudy (anonymous):

please help..

OpenStudy (imstuck):

Do you know the midpoint formula?

OpenStudy (jdoe0001):

\(\bf \textit{middle point of 2 points }\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ L&({\color{red}{ 4}}\quad ,&{\color{blue}{ -8}})\quad M&({\color{red}{ -3}}\quad ,&{\color{blue}{ 5}}) \end{array}\qquad \left(\cfrac{{\color{red}{ x_2}} + {\color{red}{ x_1}}}{2}\quad ,\quad \cfrac{{\color{blue}{ y_2}} + {\color{blue}{ y_1}}}{2} \right)\)

OpenStudy (anonymous):

yes isnt it m(Xa-Xb/2), (Ya-Yb/2)

OpenStudy (jdoe0001):

hmm,m welll ahemm ok so got the values off =) anyhow \(\bf \textit{middle point of 2 points }\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ L&({\color{red}{ 4}}\quad ,&{\color{blue}{ -3}})\quad M&({\color{red}{ -8}}\quad ,&{\color{blue}{ 5}}) \end{array}\qquad \left(\cfrac{{\color{red}{ x_2}} + {\color{red}{ x_1}}}{2}\quad ,\quad \cfrac{{\color{blue}{ y_2}} + {\color{blue}{ y_1}}}{2} \right)\)

OpenStudy (anonymous):

yes you plug it in... but how does it match any of the answer choices?

OpenStudy (imstuck):

Yes it is; just use your coordinates they gave you and find y. You'll get it if you already know the formula! Good for you!

OpenStudy (imstuck):

Your y coordinate is one of the choices they gave you. The first coordinate in the parenthesis is the x coordinate and the second coordinate is the y. They want you to identify y, even though using that formula you will also identify x.

OpenStudy (anonymous):

so would it be 3.5?

OpenStudy (imstuck):

Um...no. Let's do it together.\[(\frac{ -8+(-4) }{ 2 },\frac{ 5+3 }{ 2 })\]That gives us\[(\frac{ -12 }{ 2 },\frac{ 8 }{ 2 })\]Dividing that out we get x is -6 and y is 4. See? Choice B is 4.

OpenStudy (anonymous):

ooo ok sorry i get it i typed something wrong in my calculator. thank you so much! i understand now

OpenStudy (imstuck):

You're welcome!! Thanks for the medal!!!